OpenStudy (anonymous):

I have an exam tomorrow and need help with this problem to make sure I understand the material... Determine the y-coordinate of the critical point for f(x)=1/x+1/5-x

OpenStudy (anonymous):

Did you take the derivative?

OpenStudy (anonymous):

yea i got \[f'(x)=-1/x^2+1/(5-x)^2\]

OpenStudy (anonymous):

Looks right. So where would this function be 0?

OpenStudy (anonymous):

ummm this is where i get stuck...could i try to get x's on one side

OpenStudy (anonymous):

Well, I think he means the function will have its critical point at f'(x) = 0. Just do 1/x^2 = 1/(5-x)^2, got it?

OpenStudy (anonymous):

then would you cross multiply...???

OpenStudy (anonymous):

See, if 1/x^2 = 1/(5-x)^2, then x^2 = (5-x)^2. (x^2)^1/2 = ((5-x)^2)^1/2 |x| = |5-x| x = 5/2 or -5 = 0... since -5 is not 0, x = 5/2.

OpenStudy (anonymous):

ok i see